Find an equation of a line parallel to y=3 x+3 and passing through the point (5, 2)
Answer by Maths68(1474)

Find an equation of a line parallel to y=3 x+3 and passing through the point(5, 2)
Equation of the line slope-intercept form
y=mx+b
Given
Equation of the Given line
y=3x+3
Compare with the equation of line slope-intercept form
Slope of the given = m = 3
Since lines are parallel their slope will be same
Slope of the required line = m = 3
Point (5, 2)
We have a point and slope we can get easily the require line equation.
Use Equation of the line point-slope form
m = y2-y1/x2-x1
Put the values in above equation, we have
3=(y-2)/(x-5)
3(x-5)=(y-2)
3x-15=y-2
3x-y=-2+15
3x-y=13
-y=-3x+13
Multiply by -1 both sides of above equation
y=3x-13
Above equation is the equation of the required line